OpenStudy (anonymous):

i need tips on algabra . how do i know which number to add to each side ? for example , when i have the prob 7x-56=3x+12 then which number do i put to each side? whats the next step?

OpenStudy (anonymous):

Suppose you isolate x terms to the left side, you will add 56 to move to right side!

OpenStudy (anonymous):

sooo it looks like tho , i always use the number on the left side of the = sign . is that right?

OpenStudy (anonymous):

It's good habit to ease your confusion :)

OpenStudy (anonymous):

so id subtract 56 from 12?

OpenStudy (anonymous):

In other way, switch all variable terms to left sides: 7x - 3x switch the rest to right side: 12 + 56

OpenStudy (anonymous):

so 7x-44=3x

OpenStudy (anonymous):

Put them together: 4x = 68 -> x = 68/4 = ...

OpenStudy (anonymous):

aw thats confusing becuase i dont think we ido it like that in class

OpenStudy (anonymous):

Do whatever way that you're comfortable as long as you find solution x = ...

OpenStudy (anonymous):

well okay , but if i get 7x-44=3x next would that be right

OpenStudy (anonymous):

switch all variable terms to left sides: 7x - 3x = 4x Where do you get this 7x-44=3x ??? :O

OpenStudy (anonymous):